# Spectrum Efficiency and Processing Latency Trade-offs in Panel-Based LIS

### v1

@@ IV-B Latency Results
In Fig. 3 the latency for a system with $N=16$ and $P=128$ is shown for both ZF and MRC. When the number of UEs is relatively small, the latency distribution of different components are similar, with almost half being due to the frame structure, roughly a third due to the Ethernet-based fronthaul and the rest caused by the processing. However, when the amount of UEs grows, the two approaches lead to completely different latency ratios. For MRC, increasing the amount of UEss does not affect the latency much, as the needed computations only grow linearly with $K$ and are quite simple to execute. For ZF though, the processing latency is completely dominant for a large amount of UEs as the local processing grows quadratically with $K$ and the central processing grows cubically with $K$.
To decrease the latency of a distributed massive multiple-input multiple-output (DMIMO) system using ZF and which is designed to support a large amount of UEs, further optimizations clearly need to be done. A simple solution would be to further optimize the processing, though that would likely lead to higher resource and energy costs. Another interesting avenue that could be investigated is the potential of panel segmentation, where only a subset of panels would serve certain UEs. This would lead to smaller matrix sizes and therefore simplifying the processing, but it would cause new unique problems.
@@ V Latency Versus Spectral Efficiency
One of the main performance metrics in any wireless communication scenario is the data rate supported by the system. In the considered setting, we are specially interested in the data rates experienced by the different UEs when transmitting their UL data, and how they interplay with the arrangement of the LIS into panels, as well as with the resulting physical layer latency. To this end, we consider an industrial indoor scenario, where the channel matrix follows the Rician fading model [7]
@@
where $K_{\mathrm{F}}$ is the Rician factor, $\boldsymbol{H}_{\mathrm{LoS}}$ is the line of sight (LoS) channel between the UEs and the whole LIS, and $\boldsymbol{H}_{\mathrm{R}}$ follows the IID Rayleigh fading model where the average power for each entry equals the power of the respective LoS contribution in $\boldsymbol{H}_{\mathrm{LoS}}$. The parameters of the scenario are summarized in Table II. Note that, since we consider LIS scenarios with a large number of panels deployed throughout compact indoor regions, we should take into account near-field effects when computing the LoS contributions [12]. Specifically, we consider that only the perpendicular incident power is absorbed by each antenna element, and we construct the array steering vectors using the spherical propagation model. Moreover, we normalize the channel power such that the SNR value from Table II corresponds to the average receive SNR considering the whole LIS, where the average is taken over the whole set of simulated scenarios being compared.
@@
Room width $20$ m
We begin by analyzing the effect of distributing a fixed number of antennas into a varying number of LIS-panels. Specifically, we consider deploying $M=1024$ antennas into equidistant panels along the two walls extending through the length of the room. Note that, in all cases (except when all the antennas are co-located), each panel is facing another panel. The number of panels is incremented from $P=1$ to $P=256$ by powers of 2, while for each $P$ value the respective panels are either squared or have twice as many elements along their width than along their height. There are $K=128$ UEs transmitting equal power, and these are randomly located throughout a plane at a minimum distance of $10\lambda$ from the walls. The LIS-panels extend from the UEs plane towards the ceiling of the room. An example scenario is given in Fig. 4. All the spectral efficiency results are averaged over $10^{2}$ realizations of the uniform UE placements, each including $10^{3}$ realizations of the Rayleigh fading channel component. Fig. 5 plots the spectral efficiencies with respect to the number of panels for the described scenarios. We can see that distributing the antennas throughout the scenario into smaller LIS-panels favors user fairness, reducing potential outages for both ZF and MRC processing. We believe that this is a strong motivation for favoring more distributed antenna deployments since it increases the overall reliability. On the other hand, as seen in the previous section, in dense user deployments the latency values for the ZF case are dominated by the central processing required to perform matrix inversion, so the most distributed implementation has marginal latency increase over having all the antennas co-located in one panel. In the MRC case, the fronthaul latency associated to sharing the data sequentially between panels still has noticeable impact, leading to a latency-performance trade-off which should be further analyzed under different settings. Nevertheless, there is a major latency reduction with respect to ZF, at the cost of some degradation in average user spectral efficiency.
@@
We now consider the scaling of the spectral efficiencies when deploying varying number $P$ of equally sized square panels with $N=16$ antennas. Fig. 6 shows the average user spectral efficiency with respect to the number of panels for the indoor scenario previously described. Aside from the fixed panel size, the same scenario considerations apply as for the results in Fig. 6. As one may expect, the average user spectral efficiency increases logarithmically with the number of panels for high enough $P$. This may be justified by the extra array gain associated to having more panels receiving the transmitted power. We have also included for comparison the latency values for each panel deployment. Interestingly, the spectral efficiency achieved by MRC is comparable to that achieved by ZF with half the LIS-panels, which still requires a latency close to $30$ times higher. Thus, we can conclude that, in the considered scenario, MRC is much more effective at exploiting the trade-off between latency and spectral efficiency.
@@ VI Conclusion
We have studied the latency in a panel-based LIS indoor scenario with different panel deployments. We presented algorithms for performing decentralized MRC and ZF processing without any loss with respect to the centralized counterpart. We also computed latency values based on real implementations for the different schemes, which show that, as the number of users grows, the ZF latency is quickly dominated by the central processing required to invert the channels. Finally, we compared the spectral efficiency achieved with different deployments, where we observed that having antennas more distributed increases the system reliability, and comes at essentially no latency cost when using ZF. However, MRC seems to attain a better trade-off between latency and performance.

### v3

@@ IV-B Latency Results
In Fig. 3 the latency (absolute and relative) for a system with $N=16$ and $P=128$ is shown for ZF, MMSE and MRC. When the number of UEs is small, latency distributions are similar, with almost half of the latency due to the frame structure, roughly a third due to the Ethernet-based fronthaul, and the rest caused by the processing. However, when the amount of UEs grows, two completely different latency ratios appear. For MRC, increasing the amount of UEs does not affect the latency much, as the needed computations only grow linearly with $K$ and are simple to execute. For both ZF and MMSE though, central processing latency dominates at large numbers of UEs. While local processing grows quadratically with $K$, the central processing grows cubically with $K$.
To decrease the latency of a distributed MIMO system, using ZF or MMSE, which is designed to support a large number of of UEs, further optimizations are clearly needed. A simple solution would be to further optimize the central processing latency, specifically the matrix inversion, though that would likely lead to higher resource and energy costs. Another interesting avenue is the potential of panel segmentation, where only a subset of panels serve certain UEs. This would lead to smaller matrix sizes and, therefore, simplifying the processing, but it would, in turn, most likely cause new problems.
@@ V Latency Versus Spectral Efficiency
One of the main performance metrics in any wireless communication scenario is the data rate supported by the system. For the considered setting, we are specially interested in the data rates experienced by the different UEs in the UL, how these interplay with the arrangement of the LIS into panels, and the resulting physical layer latency. To this end, we consider an industrial indoor scenario, where the channel matrix
@@
is assumed to follow a Rician fading model [7], with $K_{\mathrm{F}}$ being the Rician factor, $\boldsymbol{H}_{\mathrm{LoS}}$ the line of sight (LoS) channel between the UEs and the entire LIS, and $\boldsymbol{H}_{\mathrm{R}}$ an IID Rayleigh fading model where average power for each entry equals the power of the respective LoS contribution in $\boldsymbol{H}_{\mathrm{LoS}}$. The parameters of the scenario are summarized in Table II. Note that, since we consider LIS scenarios with a large number of antenna elements deployed throughout compact indoor regions, we should take into account near-field effects when computing the LoS contributions [11]. Specifically, we consider that only the perpendicular incident power is absorbed by each antenna element, and we construct the array steering vectors using the spherical propagation model. Moreover, we normalize the channel power such that the SNR value from Table II corresponds to the average receive SNR considering the whole LIS, where the average is taken over the whole set of simulated scenarios being compared.
@@
Room width 2020 m
We begin by analyzing the effect of distributing a fixed number of antennas into a varying number of LIS-panels. Specifically, we consider deploying $M=1024$ antennas into equidistant panels along the two walls extending through the length of the room. Note that, in all cases (except when all the antennas are co-located), each panel is facing another panel. The number of panels is incremented from $P=1$ to $P=256$ by powers of 2, while for each $P$ value the respective panels are either square or have twice as many elements along their width than along their height. There are $K=128$ UEs transmitting equal power, and these are randomly located on a plane at a minimum distance of $10\lambda$ from the walls. The LIS-panels extend from the UEs plane towards the ceiling of the room. Two example scenarios are shown in Fig. 4. All spectral efficiency results are averaged over $10^{2}$ realizations of the uniform UE placements, each including $10^{3}$ realizations of the Rayleigh fading channel component. Fig. 5 plots the spectral efficiencies with respect to the number of panels for the described scenarios. We can see that distributing the antennas throughout the scenario into more smaller LIS-panels favors user fairness, reducing potential outages for all processing schemes, while it also allows for a slight increase in average user rate. We believe that this is a strong motivation for favoring more distributed antenna deployments since it increases the overall system reliability. At the same time, as seen in the previous section, in dense user deployments the latency values for the ZF and MMSE cases are dominated by the central processing required to perform matrix inversion, so the most distributed implementation only has marginal latency increase over having all the antennas co-located in one panel. In the MRC case, the fronthaul latency associated to sharing the data sequentially between panels still has noticeable impact, leading to a latency-performance trade-off which should be further analyzed under different settings. Nevertheless, there is a major latency reduction with respect to ZF and MMSE, at the cost of some degradation in average user spectral efficiency.
@@
We now consider the scaling of the spectral efficiencies when deploying varying numbers, $P$, of equally sized square panels with $N=16$ antennas each. Fig. 6 shows the average user spectral efficiency with respect to the number of panels for the indoor scenario previously described. Aside from the fixed panel size, the same scenario considerations apply as for the results in Fig. 5. For large enough $P$, ZF and MMSE exhibit similar performance, hinting that such scenario is interference limited, while the average user spectral efficiency increases logarithmically with the number of panels, which is justified by the extra array gain associated to having more panels receiving the transmitted power. Fig. 6 also includes the associated latency values for each panel deployment. Interestingly, the spectral efficiency achieved by MRC is comparable to that achieved by ZF and MMSE with a quarter of the LIS-panels, which still requires a latency close to $30$ times higher. Thus, we conclude that, in the considered scenario, MRC is much more effective at exploiting the trade-off between latency and spectral efficiency.
@@ VI Conclusion
We have studied the latency in a panel-based LIS indoor scenario with different panel deployments. We presented algorithms for performing decentralized MRC, ZF and MMSE processing without any loss compared to the centralized counterpart. We also compute latency values based on real implementations for the different schemes, which show that, as the number of users grows, the ZF and MMSE latency is quickly dominated by the central processing required to invert the channel Gramian. Finally, we compared the spectral efficiency achieved with different deployments, where we observed that having antennas more distributed increases the system reliability, and comes at essentially no extra latency cost when using ZF or MMSE. However, MRC seems to attain a better trade-off between latency and performance.